instructions: evaluate the discriminant of each equation. determine how many real solutions each equation has.
1. 16x^2-8x+1=0

Answer by stanbon(75887)

evaluate the discriminant of each equation. determine how many real solutions each equation has.
16x^2-8x+1 = 0
------
b^2-4ac = (-8)^2 - 4*16*1 = 64 - 64 = 0
-----------
Ans: 2 equal Real solutions.
=====================================
Cheers,
Stan H.